OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of 3/4

OpenStudy (anonymous):

y + 1 = 3/4 (x - 2)

OpenStudy (anonymous):

y = 3/4x - 3/2 -2/2 y = 3/4x -5/2

OpenStudy (anonymous):

You can use point-slope form here: \[(y - y _{1}) = m (x - x _{1})\] Substitute your slope into m. And your y-coordinate for y1 and your x-coordinate for x1. Then solve for y.

OpenStudy (anonymous):

_ x _ _ y _ _ _ thats how i have to put it in

OpenStudy (anonymous):

Using point slope the equation looks like this: y + 1 = 3/4 (x - 2) y + 1 = 3/4x - 6/4 y = 3/4 x - 6/4 - 4/4 y = 3/4 x - 5/2